# Turbulent gas-rich discs at high redshift: origin of thick stellar discs through 3D ‘baryon sloshing’

### v1

@@ 2.2 Reference frames
An idealised, multi-component galaxy system is made up of several distinct contributions to the gravitational potential. Since the initial galaxy is set up as an equilibrium figure using agama, this inertial reference frame conserves momentum in all components, for both gas, stars and dark matter (Vasiliev, 2019). In N-body simulations, the different mass components can become separated, depending on the nature and dynamics of the dark matter halo. In one of their examples, Joshi & Widrow (2024) plot the separate trajectories for the centre of mass of the halo, bulge and disc. Once the baryonic component starts to slosh, since each subsystem has its own internal structure, the restoring force and the drag imposed by dynamical friction can be very different.
As the gas fraction goes up, there is an increasing amount of noise in the system, particularly from the large-scale turbulent energy maintained by the star formation. As we show, this causes a strong sloshing of the baryons within the total gravitational potential well. We find it useful and revealing to define two distinct centers that are tied to two different reference frames, with the first being defined by the minimum of the baryonic gravitational potential:
@@
The second reference frame defines the minimum of the total potential (dominated by dark matter):
@@
with all partial second derivatives for $\Phi_{\rm tot}$ being positive definite. This centre of the total potential ($x_{\rm tot}$, $y_{\rm tot}$, $z_{\rm tot}$) is unseen in observations and, typically, in sloshing discs, only inferred from the disc kinematics. In practice, the minima defined in Eqs. 3 and 4 are identified after convolving the gravitational field at each timestep with a window function defined by 50 nearest-neighbour particles. Given that the central baryons dominate over dark matter, we anticipate that both minima have similar sloshing amplitudes with respect to an inertial frame.
@@ 3.1 Unbinding the disc
In Fig. 3, we show how the radial profile of the simulated disc’s gravitational potential $\Phi_{\rm bary}(R,t)$ evolves with time $t$ as a function of the disc gas fraction $f_{\rm gas}$. For the ‘no halo accretion’ models, there is only a 15--20% change for $f_{\rm gas}<0.5$. In the halo accretion models, there is substantial accretion to the centre at low $f_{\rm gas}$, but the overall radial profile is relatively unchanged for $f_{\rm gas}<0.5$. At higher $f_{\rm gas}$, the gravitational potential decreases rapidly over the first 500 Myr for both accretion cases. Up to half of the gravitational potential is lost due to gas launched into the halo by strong disc-wide winds. Over the next 1.5 Gyr, there is a slow re-accretion to the inner disc, but most of it is recirculated to the outer disc through recycling in the galactic halo. The slow spin of the halo ensures the gas acquires angular momentum. We explore the physics of this process in a later paper.
A great deal of energy is needed to ‘unbind’ the disc in this way, but the feedback requirements through active star formation are consistent with high-redshift ALMA and JWST observations. We can obtain a rough estimate from the following formula:
@@
Alternatively, the timescale $\Delta\tau$ needed to partially unbind (loosen) the disc is essentially the time it takes to lose half of its mass or, equivalently, its surface density. Thus
@@
where $\dot{\Sigma}=\epsilon\dot{\mu_{*}}$, and where $\epsilon$ ($=0.1$) is the coupling efficiency of the feedback energy and radiation with the gas (Egorov et al., 2023, see their Sec. 5.2). Here, $\Delta\tau$ is the duration of the starburst,
@@
The term $\dot{\mu}_{\star}$ is the SFR surface density; this is the quantity shown in Fig. 2 averaged over annular bins at each radius. Therefore,
@@
This is about $0.3\pm 0.2$ M kpc-2 yr-1 at $R\approx 8$ kpc where the disc is more reactive and easily unbound, consistent with Fig. 2, where sustained bursts over 100 Myr are a regular occurrence. The simulated bursts can be much higher but these are occurring in compact regions and on shorter timescales than $\Delta\tau$. The inferred time-averaged rates are sufficiently impulsive to (partially) unbind the disc, particularly at large galactocentric radius. Here, we have taken the $R_{\rm disc}$ ($=1.7\pm 0.3$ kpc) and $\Sigma_{0}$ normalizations from our simulated models (see below). At these early times, the most powerful starbursts create of order $3\times 10^{9}$ stars in a $\Delta\tau$ interval across the inner kiloparsec of the disc.
@@ 3.2 Stochastic fluctuations & increasing entropy
There is an interesting precedent for the disc inflation process. In the context of dwarf galaxies, Pontzen & Governato (2012) show how fluctuations in the central gravitational potential $\Phi_{0}$ irreversibly transfer energy into collisionless particles, thus removing any dark matter ‘cusp’ predicted in N-body simulations (Navarro et al., 1996). They consider blowout-recollapse cycles driven by a central starburst (see also Ogiya & Mori, 2011; Nipoti & Binney, 2015). By analogy, consider a star that is oscillating vertically with orbital frequency $\nu$ under simple harmonic motion
@@
where $z_{\rm disc}$ is the disc’s vertical scale length. (The formal definition is given below.) The key point is that a prompt change (i.e. faster than the orbital period) of $\nu$ at a fixed radius due to a sudden blowout and recollapse is now associated with a different amplitude $a_{1}$ and phase $\psi_{1}$ along the star’s orbit compared to the original amplitude $a_{0}$ and phase $\psi_{0}$ (cf. Yu et al., 2021). According to Pontzen & Governato (2012), the change in energy has an explicit dependence on the orbital phase, such that
@@
Figure 6: Nexus simulation for an $f_{\rm disc}=0.5$ stellar disc with no cold gas component ($f_{\rm gas}=0$%). The top left and bottom left panels are the $x-y$ and $x-z$ projections respectively at a fixed timestep (the boxed information concerns the disc asymmetry and concentration, a topic discussed in our companion paper, Bland-Hawthorn et al. 2025). On the right, a small amount of ‘baryon sloshing’ ($\lesssim 50$ pc) is expected due to the discreteness of the disc, as demonstrated. The colour coding indicates the passage of time, as indicated. Once again, top right and bottom right panels are the $x-y$ and $x-z$ projections respectively. The baryon sloshing is defined in terms of the separation of the minimum of the total gravitational potential ($\Phi_{\rm tot}(0,0)$ indicated by the cross at the origin) and the baryon potential ($\Phi_{\rm bary}(0,0)$).
Figure 7: Nexus simulations in the presence of halo accretion for an $f_{\rm disc}=0.5$ stellar disc for different values of $f_{\rm gas}$, as indicated. Left panel: $x-y$ (left column) and $x-z$ (right column) projections of the sloshing action. Note that the physical scale of each plot is growing with increasing $f_{\rm gas}$. Right panel: $x-y$ (left column) and $x-z$ (right column) projections of the sloshing action. Note that the physical scale of each plot is fixed for all $f_{\rm gas}$. The baryon sloshing is defined in terms of the separation of the minimum of the total gravitational potential $\Phi_{\rm tot}(0,0,0)$ (indicated by a cross) and the baryon gravitational potential, $\Phi_{\rm bary}(0,0,0)$. The colour coding indicates the passage of time, as indicated by the RHS bar. Internal to each plot, the horizontal and vertical scales are always matched.
@@ 3.3 Sloshing behaviour
Figure 8: The 2D slosh projections in Fig. 7 can be rendered in 3D to reveal the erratic movement. The tracks all refer to the $f_{\rm gas}=100\%$ halo accretion case: (Left) Motion of $\Phi_{\rm tot}(0,0,0)$ with respect to the inertial reference frame; (Middle) Motion of $\Phi_{\rm bary}(0,0,0)$ with respect to the inertial reference frame; (Right) Motion of $\Phi_{\rm bary}(0,0,0)$ with respect to $\Phi_{\rm tot}(0,0,0)$. The matched $xyz$ axes are in units of kpc in a comoving frame. Evidently, for the baryons, there are frequent major reversals over the 2 Gyr simulation. The colour coding indicates the passage of time, from blue to red over this timespan.
@@
The baryon sloshing is defined in terms of the separation of the minimum of the total gravitational potential ($\Phi_{\rm 0,tot}$ indicated by the cross) and the baryon potential ($\Phi_{\rm 0,bary}$) defined in equations 4 and 3 respectively. In Fig. 7, we repeat the exercise carried out in Fig. 6 to show the sloshing behaviour for $f_{\rm gas}=20,40,60,80,100$%. The “post-processed” movies at our website http://www.physics.usyd.edu.au/turbo_discs/ clearly show the baryon sloshing in the inertial frame. See, for example, the $x-y$ projection of the newly formed stars (left middle panel) in fd50_fg100_ac; the emerging stellar bar soon drifts away from the origin and follows a complex track with multiple loops.
@@
where $M_{\rm halo}$ is the total halo mass (dark matter $+$ hot coronal gas), and ${\bm{V}}_{\rm disc}$ is the disc’s velocity with respect to the total gravitational potential. Thus, as expected, since $\nu>\kappa$, the sloshing perturbations for all simulations is more compressed along the vertical axis when compared with in-plane motions.
@@ 4 The origins of sloshing
The dynamics of stellar discs is a remarkably complex process, with only restricted analytic cases that are well understood (Binney, 2025). So to what extent can we trust N-body simulations to trace the growth of instabilities and non-linear perturbations realistically? One of the most powerful analytic theories to address this issue is the (local) shearing-sheet model presented by Julian & Toomre (1966) for a razor-thin 2D disc. Here, a perturbation in the distribution function (DF) can be followed in both density space and velocity space. A local disturbance generates a leading wake that is “swing-amplified” into a trailing wake. These density perturbations grow as a consequence of gravity, shear and rotation (Toomre, 1964), but eventually dissolve unless sustained by a perturbing (e.g. external) force. Even when the density enhancement has disappeared, its influence is seen in velocity space long afterwards but eventually fades from view due to phase-mixing.
@@
In the shearing sheet approximation, the swing-amplified modes are analysed in a frozen, co-rotating frame where both the global and local rotation (due to shear and vorticity) are removed. The model treats stars that are on near-circular orbits, although corrections can be made for mildly non-circular orbits (Binney, 2020). In such a frame, all curvature is removed and spiral features can be treated as time-dependent plane waves, typically via a local density (or surface density) -- gravitational potential pair of coupled functions.
4.2 Jolting the disc -- the importance of self-gravity
@@
So how does the shearing sheet respond to a powerful jolt? This line of enquiry is revealing because it shows that an impulse experienced by a stellar disc is amplified by its self-gravity for all $Q_{\star}$ $(=\kappa\sigma_{R}/3.36G\Sigma_{0})$ values relevant to Nexus galaxies. Here, $\Sigma_{0}$ is the local undisturbed surface density, $\kappa$ is the epicyclic frequency and $\sigma_{R}$ is the radial velocity dispersion. $Q_{\star}$ measures the local axisymmetric stability in the stellar disc; non-axisymmetric growing modes are still allowed. The amplifying effect is strongest at intermediate wavelengths (see below), becoming weaker at long wavelengths, while short wavelengths are suppressed below a critical wavelength $\lambda_{\rm crit}=2\pi/k_{\rm crit}$, for which $k_{\rm crit}=\kappa^{2}/2\pi G\Sigma_{0}$. The characteristic epicyclic radius is roughly $\sigma_{R}/\kappa$. These physical dimensions are on kiloparsec scales, e.g. $k_{\rm crit}\approx(2\pm 1\rm\;kpc)^{-1}$, such that $\lambda_{\rm crit}\sim 10$ kpc near the Solar circle.
@@
In the shearing-sheet approximation (Julian & Toomre, 1966), the coordinate system is centred on a spiralling wave’s corotation radius. In Galactic cylindrical coordinates ($R,\phi,z$), individual stars have velocities ${\bm{V}}$ $=$ ($V_{R}$, $V_{\phi}$, $V_{z}$) and oscillation frequencies ${\bm{\Omega}}$ $=$ ($\Omega_{R},\Omega_{\phi},\Omega_{z}$) $=$ ($\kappa,\omega,\nu$). Following Binney (2020), we introduce a cartesian wave vector ${\bm{k}}$ that corotates around a point defined by the radius vector $\bm{R}$ such that
@@
where $k_{y}$ is the tangential wavenumber, and ${\cal A}=-\frac{1}{2}\frac{d\Omega_{\phi}}{d\ln r}$ is Oort’s constant that quantifies the shear rate. The wave vector ensures that the perturbation’s wave crests track from leading ($t<t_{0},k_{x}<0$) to radial ($t=t_{0},k_{x}=0$) to trailing waves ($t>t_{0},k_{x}>0$), where the amplitude peaks near $t=t_{0}$. A suitable potential -- surface density pair is
@@
where $\bm{x}$ describes the $x-y$ plane, and $\Sigma_{0}$ and $\Sigma_{1}$ are the undisturbed and wave-disturbed density fields.
@@
In Fig. 9, we show the response of a razor-thin 2D disc to a strong local impulse for different values of the $Q$ parameter, from $Q=1$ (marginally stable, high self-gravity) to $Q=1.4$ (most stable, low self-gravity). [^5] We acknowledge and thank James Binney for providing the code for this calculation (see Binney, 2020). We initiated the impulse at a time $t=t_{i}$ before $t=t_{0}$; the perturbed distribution function is $f_{1}(t_{i})=0$ at the start. The impulse is such that the surface density from the perturbation is $\tilde{\Sigma}_{e}(t,t_{i})=\delta(t-t_{i})/\kappa$. Time is needed for the perturbation to build, but if that time is set too far in the past, the disturbance can phase-mix away before it has a chance to build. From the so-called JT equation, the renormalized surface-density evolution is given by
[…]
@@
In the Nexus simulations, away from the galaxy’s centre, $Q$ is in the range $1.2-1.5$, leading to swing-amplification factors $A$ as high as $10-60$ compared to discs without self-gravity (cf. Aumer et al., 2016). Binney (2020) showed that the biggest amplification occurs around $\lambda/\lambda_{\rm crit}\approx 2\pm 1$ (see also Toomre, 1981). Thus, it is entirely reasonable that a succession of disc jolts leads to a strong reaction from the large-scale stellar disc on tens of kiloparsec scales. Once the momentum-injecting starbursts die down, the disc ceases to slosh, as observed.
Figure 10: Two different perspectives of the same Brownian motion track in a moderately flattened, spherical gravitational potential. The colour scheme once again shows the evolution over 2 Gyr (blue $\rightarrow$ red). The track is similar in character to the baryon sloshing tracks in Figs. 7 and 8.

### v2

@@ 2.3 Reference frames
An idealised, multi-component galaxy system is made up of several distinct contributions to the gravitational potential. Since the initial galaxy is set up as an equilibrium figure using agama (Vasiliev, 2019), this inertial reference frame conserves momentum in all components, for both gas, stars and dark matter. In N-body simulations, the different mass components can become separated, depending on the nature and dynamics of the dark matter halo. In one of their examples, Joshi & Widrow (2024) plot the separate trajectories for the centre of mass of the halo, bulge and disc. Once the baryonic component starts to slosh, since each subsystem has its own internal structure, the restoring force and the drag imposed by dynamical friction can be very different.
As the gas fraction increases, the system experiences stronger gravitational potential fluctuations, primarily driven by large-scale turbulent energy sustained by ongoing star formation. In particular, momentum recoil from sporadic, galaxy-wide starbursts adds significantly to these background fluctuations. As we demonstrate, the cumulative effect is a pronounced sloshing of the baryonic matter within the overall gravitational potential well. We find it useful and revealing to define two distinct centers that are tied to two different reference frames, with the first being defined by the minimum of the baryonic gravitational potential:
@@
The second reference frame defines the minimum of the total potential:
@@
with all partial second derivatives for $\Phi_{\rm tot}$ being positive definite. This centre of the total potential ($x_{\rm tot}$, $y_{\rm tot}$, $z_{\rm tot}$) is not traced by light, typically, in sloshing discs, only inferred from the disc kinematics. In practice, the minima defined in Eqs. 3 and 4 are identified after convolving the gravitational field at each timestep with a window function defined by 50 nearest-neighbour particles. Given that the central baryons dominate over dark matter, we anticipate that both minima have similar sloshing amplitudes with respect to an inertial frame.
@@ 3.1 Unbinding the disc
In Fig. 3, we show how the radial profile of the simulated disc’s gravitational potential $\Phi_{\rm bary}(R,t)$ evolves with time $t$ as a function of the disc gas fraction $f_{\rm gas}$. For the ‘no halo accretion’ models, there is only a 15-20% change for $f_{\rm gas}<0.5$. In the halo accretion models, there is substantial accretion to the centre at low $f_{\rm gas}$, but the overall radial profile is relatively unchanged for $f_{\rm gas}<0.5$. At higher $f_{\rm gas}$, the gravitational potential decreases rapidly over the first 500 Myr for both accretion cases. Up to half of the gravitational potential is lost due to gas launched into the halo by strong disc-wide winds. (The same process has been discussed in the context of dwarf galaxies by Zhao (2002).) Over the next 1.5 Gyr, there is a slow re-accretion to the inner disc, but most of it is recirculated to the outer disc through recycling in the galactic halo. The slow spin of the halo ensures the gas acquires angular momentum. We will explore the physics of this process in a later paper.
A great deal of energy is needed to ‘unbind’ the disc in this way, but the feedback requirements through active star formation are consistent with high-redshift ALMA and JWST observations. We can obtain a rough estimate from the following formula for the binding energy:
@@
Alternatively, the timescale $\Delta\tau$ needed to partially unbind (loosen) the disc is essentially the time it takes to lose half of its mass or, equivalently, its surface density (e.g. Hills, 1980). Thus
@@
where $\dot{\Sigma}=\epsilon\dot{\mu_{*}}$, for which $\dot{\mu}_{\star}$ is the SFR surface density, and where $\epsilon$ (=0.1=0.1) is the coupling efficiency of the feedback energy and radiation with the gas (Egorov et al., 2023, see their Sec. 5.2). Here, $\Delta\tau$ is the duration of the starburst,
@@
The SFR surface density $\dot{\mu}_{\star}$ is the quantity shown in Fig. 2 averaged over annular bins at each radius. Therefore, after rearranging Eq. 8 and inserting the simulated surface density profile $\Sigma(R)$,
@@
This is about $\dot{\mu}_{\star}=0.3\pm 0.2$ M kpc-2 yr-1 at $R\approx 8$ kpc where the disc is more reactive and easily unbound, consistent with Fig. 2, where sustained bursts over 100 Myr are a regular occurrence, hence the choice of normalization for $\Delta\tau$ in the above equation. The simulated bursts can be much higher but these are occurring in compact regions and on shorter timescales than $\Delta\tau$. The inferred time-averaged rates are sufficiently impulsive to (partially) unbind the disc, particularly at large galactocentric radius. Here, we have taken the $R_{\rm disc}$ ($1.7\pm 0.3$ kpc) and $\Sigma_{0}$ normalizations from our simulated models (see below). At these early times, the most powerful starbursts create of order $3\times 10^{9}$ stars in a $\Delta\tau$ interval across the inner kiloparsec of the disc.
@@ 3.2 Stochastic fluctuations & increasing entropy
There is an interesting precedent for the disc inflation process in the context of dwarf galaxies (Zhao, 2002). A more extensive and related literature focusses on unbinding the NFW cusp predicted in dark-matter only simulations, but never observed. For example, Pontzen & Governato (2012) show how fluctuations in the central gravitational potential $\Phi_{0}$ irreversibly transfer energy into collisionless particles, thus removing any dark matter ‘cusp’ (Navarro et al., 1996). They consider blowout-recollapse cycles driven by a central starburst (see also Ogiya & Mori, 2011; Nipoti & Binney, 2015). By analogy, consider a star that is oscillating vertically with orbital frequency $\nu$ under simple harmonic motion
@@
where $z_{\rm disc}$ is the disc’s vertical scale height. (The formal definition is given below.) The key point is that a prompt change (i.e. faster than the orbital period) of $\nu$ at a fixed radius due to a sudden blowout and recollapse is now associated with a different amplitude $a_{1}$ and phase $\psi_{1}$ along the star’s orbit compared to the original amplitude $a_{0}$ and phase $\psi_{0}$ (cf. Yu et al., 2021). According to Pontzen & Governato (2012), the change in energy has an explicit dependence on the orbital phase, such that
@@
Figure 6: Nexus simulation for an $f_{\rm disc}=0.5$ stellar disc with no cold gas component ($f_{\rm gas}=0$%). The top left and bottom left panels are the xyx-y and xzx-z slosh projections respectively at a fixed timestep. The baryon sloshing is defined in terms of the separation of the minimum of the total gravitational potential ($\Phi_{\rm tot}(0,0)$ indicated by the cross at the origin) and the baryon potential ($\Phi_{\rm bary}(0,0)$). (The boxed information concerns the disc asymmetry and concentration, a topic discussed in our companion paper, Bland-Hawthorn et al. 2025. The physical scales of the 3 circles are indicated in the lower key). The top right and bottom right panels are the xyx-y and xzx-z slosh projections respectively, but highly magnified compared to the LHS panels. A small amount of ‘baryon sloshing’ ($\lesssim 50$ pc) is expected due to the discreteness of the disc, as demonstrated. The colour coding indicates the baryon centroid shifts with the passage of time, as indicated.
Figure 7: Nexus simulations in the presence of halo accretion for an $f_{\rm disc}=0.5$ stellar disc for different values of $f_{\rm gas}$, as indicated. Left panel: xyx-y (left column) and xzx-z (right column) projections of the sloshing action. Note that the physical scale of each plot is growing with increasing $f_{\rm gas}$. Right panel: xyx-y (left column) and xzx-z (right column) projections of the sloshing action. Note that the physical scale of each plot is fixed for all $f_{\rm gas}$. The baryon sloshing is defined in terms of the separation of the minimum of the total gravitational potential $\Phi_{\rm tot}(0,0,0)$ (indicated by a cross) and the baryon gravitational potential, $\Phi_{\rm bary}(0,0,0)$. The colour coding indicates the passage of time, as indicated by the RHS bar, extending from $z\approx 3.3$ to $z\approx 1.5$. Internal to each plot, the horizontal and vertical scales are always matched.
@@ 3.3 Sloshing behaviour
Figure 8: The 2D slosh projections in Fig. 7 can be rendered in 3D to reveal the erratic movement. The tracks all refer to the $f_{\rm gas}=100\%$ halo accretion case: (Left) Motion of $\Phi_{\rm tot}(0,0,0)$ with respect to the inertial reference frame; (Middle) Motion of $\Phi_{\rm bary}(0,0,0)$ with respect to the inertial reference frame; (Right) Motion of $\Phi_{\rm bary}(0,0,0)$ with respect to $\Phi_{\rm tot}(0,0,0)$. The slosh axes are in units of kpc in a comoving frame. Evidently, for the baryons, there are frequent major reversals over the 2 Gyr simulation. The colour coding indicates the passage of time, from blue to red over this timespan.
@@
The baryon sloshing is defined in terms of the separation of the minimum of the total gravitational potential ($\Phi_{\rm 0,tot}$ indicated by the cross) and the baryon potential ($\Phi_{\rm 0,bary}$) given by equations 4 and 3 respectively. In Fig. 7, we repeat the exercise carried out in Fig. 6 to show the sloshing behaviour for $f_{\rm gas}=20,40,60,80,100$%. The “post-processed” movies at our website http://www.physics.usyd.edu.au/turbo_discs/ clearly show the baryon sloshing in the inertial frame. See, for example, the xyx-y projection of the newly formed stars (left middle panel) in fd50_fg100_ac; the emerging stellar bar soon drifts away from the origin and follows a complex track with multiple loops.
@@
where $M_{\rm halo}$ is the total halo mass (dark matter ++ hot coronal gas), and ${\bm{V}}_{\rm disc}$ is the disc’s velocity with respect to the total gravitational potential. Thus, as expected, since $\nu>\kappa$, the sloshing perturbations for all simulations is more compressed along the vertical axis when compared with in-plane motions.
@@ 4 The origins of sloshing
The dynamics of stellar discs is a remarkably complex process, with only restricted analytic cases that are well understood (Binney, 2025; Hamilton et al., 2025). So to what extent can we trust N-body simulations to trace the growth of instabilities and non-linear perturbations realistically? One of the most powerful analytic theories to address this issue is the (local) shearing-sheet model presented by Julian & Toomre (1966) for a razor-thin 2D disc. Here, a perturbation in the distribution function (DF) can be followed in both density space and velocity space. A local disturbance generates a leading wake that is “swing-amplified” into a trailing wake. These density perturbations grow as a consequence of gravity, shear and rotation (Toomre, 1964), but eventually dissolve unless sustained by a perturbing (e.g. external) force. Even when the density enhancement has disappeared, its influence is seen in velocity space long afterwards but eventually fades from view due to phase-mixing.
@@
In the shearing sheet approximation, the swing-amplified modes are analysed in a frozen, co-rotating frame where both the global and local rotation (due to shear and vorticity) are removed. The model treats stars that are on near-circular orbits, although corrections can be made for mildly non-circular orbits (Binney, 2020). In such a frame, all curvature is removed and spiral features can be treated as time-dependent plane waves, typically via a local density (or surface density) - gravitational potential pair of coupled functions.
4.2 Jolting the disc - the importance of self-gravity
@@
So how does the shearing sheet respond to a powerful jolt? This line of enquiry is revealing because it shows that an impulse experienced by a stellar disc is amplified by its self-gravity for all $Q_{\star}$ $(=\kappa\sigma_{R}/3.36G\Sigma_{0})$ values relevant to Nexus galaxies. Here, $\Sigma_{0}$ is the local undisturbed surface density, $\kappa$ is the epicyclic frequency and $\sigma_{R}$ is the radial velocity dispersion. $Q_{\star}$ is a dimensionless quantity that measures the local axisymmetric stability in the stellar disc against collapse or fragmentation; non-axisymmetric growing modes are still allowed. The amplifying effect is strongest at intermediate wavelengths (see below), becoming weaker at long wavelengths, while short wavelengths are suppressed below a critical wavelength $\lambda_{\rm crit}=2\pi/k_{\rm crit}$, for which $k_{\rm crit}=\kappa^{2}/2\pi G\Sigma_{0}$. The characteristic epicyclic radius is roughly $\sigma_{R}/\kappa$. These physical dimensions are on kiloparsec scales, e.g. $k_{\rm crit}\approx(2\pm 1\rm\;kpc)^{-1}$, such that for example $\lambda_{\rm crit}\sim 10$ kpc near the Solar circle today.
@@
In the shearing-sheet approximation (Julian & Toomre, 1966), the coordinate system is centred on a spiralling wave’s corotation radius. In Galactic cylindrical coordinates ($R,\phi,z$), individual stars have velocities $\bm{V}$} == ($V_{R}$, $V_{\phi}$, $V_{z}$) and oscillation frequencies $\bm{\Omega}$} == ($\Omega_{R},\Omega_{\phi},\Omega_{z}$) == ($\kappa,\omega,\nu$). Following Binney (2020), we introduce a cartesian wave vector $\bm{k}$} that corotates around a point defined by the radius vector $\bm{R}$ such that
@@
where $k_{y}$ is the tangential wavenumber, and ${\cal A}=-\frac{1}{2}\frac{d\Omega_{\phi}}{d\ln r}$ is Oort’s constant that quantifies the shear rate. The wave vector ensures that the perturbation’s wave crests track from leading ($t<t_{0},k_{x}<0$) to radial ($t=t_{0},k_{x}=0$) to trailing waves ($t>t_{0},k_{x}>0$), where the amplitude peaks near $t=t_{0}$. A suitable potential - surface density pair is
@@
where $\bm{x}$ describes the xyx-y plane, and $\Sigma_{0}$ and $\Sigma_{1}$ are the undisturbed and wave-disturbed density fields.
@@
In Fig. 9, we show the response of a razor-thin 2D disc to a strong local impulse for different values of the $Q$ parameter, from $Q=1$ (marginally stable, high self-gravity) to $Q=1.4$ (most stable, low self-gravity). We initiated the impulse at a time $t=t_{i}$ before $t=t_{0}$; the perturbed distribution function is $f_{1}(t_{i})=0$ at the start. The impulse is such that the surface density from the perturbation is $\tilde{\Sigma}_{e}(t,t_{i})=\delta(t-t_{i})/\kappa$. Time is needed for the perturbation to build, but if that time is set too far in the past, the disturbance can phase-mix away before it has a chance to build. From the so-called JT equation, the renormalized surface-density evolution is given by
$\tilde{\Sigma}_{1}(t)$ $=$ $\int_{t_{i}}^{t}{\cal K}(t,t^{\prime})(\tilde{\Sigma}_{1}(t^{\prime})+\tilde{\Sigma}_{e}(t^{\prime}))\;dt^{\prime}$ (22)
$=$ ${\cal K}(t,t_{i})+\int_{t_{i}}^{t}{\cal K}(t,t^{\prime})\tilde{\Sigma}_{1}(t^{\prime})\;dt^{\prime}\,.$ (23)
@@
In the Nexus simulations, away from the galaxy’s centre, $Q$ is in the range 1.21.51.2-1.5, leading to swing-amplification factors $A$ as high as 106010-60 compared to discs without self-gravity (cf. Aumer et al., 2016). Binney (2020) showed that the biggest amplification occurs around $\lambda/\lambda_{\rm crit}\approx 2\pm 1$ (see also Toomre, 1981). Thus, it is entirely reasonable that a succession of disc jolts leads to a strong reaction from the large-scale stellar disc (e.g. spiral arm excitation, internal disc heating) on tens of kiloparsec scales (q.v. Hamilton et al., 2025). Once the momentum-injecting starbursts die down, the disc ceases to slosh, as observed in the simulations.
Figure 10: Two different perspectives of the same Brownian motion track in a moderately flattened, spherical gravitational potential. The slosh coordinates (in kpc) are equivalent to those presented in Fig. 7. The colour scheme once again shows the evolution over 2 Gyr (blue $\rightarrow$ red). The track is similar in character to the baryon sloshing tracks in Figs. 7 and 8.